# Markov Perfect Equilibria in

### v1

@@ 3.1 Existence
$\begin{cases}v^{\beta}(t,x,z)=\underset{a\in A(h,x)}{\inf}\,H(x,z,E(x,z,v^{\beta}(t+h),\beta(t)),a),\quad 0\leq t\leq T,\\ \alpha(t,x,z)=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\ H(x,z,E(x,z,v^{\beta}(t+h),\beta(t)),a),\quad 0\leq t<T,\\ v^{\beta}(T,x,z)=g(x,z),\end{cases}$
by backward recursion. Then set $\Phi(\beta):=\alpha$. Using Berge’s maximum theorem, it is easy to see that $\Phi$ is a continuous mapping. Since $\mathscr{A}_{N}$ is a convex compact set, Brouwer’s fixed-point theorem guarantees the existence of a fixed point. By Remark 2.6 any fixed point corresponds to a symmetric Markov perfect equilibrium. ∎
@@ 3.2 Uniqueness
where $\{\xi^{i}\}$ are independent $\mathrm{Uniform}(0,1)$ variables. Now fix $j\in{\mathcal{X}}$ and $(x,z)\in\mathscr{X}_{N}$, let $\{\xi^{i}_{y}\}$ be a collection of independent $\mathrm{Uniform}(0,1)$ random variables, and write
$Z^{\prime}_{j}(x,z,\alpha)-Z^{\prime}_{j}(x,z,\tilde{\alpha})$ $=\sum_{y\in{\mathcal{X}}}(Z^{\prime}_{j}(x,y,z,\alpha)-Z^{\prime}_{j}(x,y,z,\tilde{\alpha}))$
@@
Taking expectations,
$\mathbb{E}|Z^{\prime}(x,z,\alpha)-Z^{\prime}(x,z,\tilde{\alpha})|_{1}$ $=\sum_{j\in{\mathcal{X}}}\mathbb{E}|Z^{\prime}_{j}(x,z,\alpha)-Z^{\prime}_{j}(x,z,\tilde{\alpha})|$
@@
$\leq\,(1+\max_{y\in{\mathcal{X}}}\mathrm{Lip}_{1}(\alpha(y,\cdot)))\,|z-\tilde{z}|_{1}.\qed$
We note the following version of the conjugate correspondence theorem, whose proof is given in Appendix A.2.
@@
Let $f:\mathbb{R}^{d}\to(-\infty,\infty]$ be a proper, closed and convex function. Let $C\subset\mathrm{dom}(f)$ be a non-empty, closed and convex set. Fix $p,q\in[1,\infty]$ with $11/p+1/q=$ 1 and assume that there exists a $\gamma>0$ such that
$f(x^{\prime})\geq f(x)+\phi\cdot(x^{\prime}-x)+\gamma|x^{\prime}-x|_{p}|x^{\prime}-x|_{q}$ (3.1)
for any $x,x^{\prime}\in C$ and $\phi\in\partial f(x)$. Set
$\psi(y)=\underset{x\in C}{\mathrm{arg\,max}}\{y\cdot x-f(x)\},\quad y\in\mathbb{R}^{d}.$
Then, $\psi$ is single-valued and is Lipschitz continuous with constant $1/\gamma$, with respect to both $|\cdot|_{p}$ and $|\cdot|_{q}$.
We are now ready to prove existence and uniqueness for one-step games. We recall the constant $\gamma$ from Assumption 2.7, $S(x)$ and $m=m(d)$ from Definition 2.2, and the constant $L_{\partial\ell}$ from Assumption 2.19.
@@
Let Assumptions 2.5 and 2.7 hold. Then, for any $\varphi:\mathscr{X}_{N}\to\mathbb{R}$, there exists a unique function $\hat{\alpha}:\mathscr{X}_{N}\to\Sigma^{d-1}$ that satisfies
@@
For $\alpha\in{\mathcal{S}}$, set $|\alpha|_{\mathcal{S}}=\max_{(x,z)\in\mathscr{X}_{N}}|\alpha|_{1}$,
@@
Now let $(x,z)\in\mathscr{X}_{N}$ be given. Define a function $f:\mathbb{R}^{d}\to[0,\infty]$ by $f(a):=\ell(x,z,a/h)h$ for $a\in[0,\infty)^{d}$ and +$\infty$+ ∞ otherwise. Set $C:=A(h,x)$. By Assumptions 2.5 and 2.7, $f$ is a proper, closed and convex function. Using Assumption 2.7 and the fact that any element $a\in C$ satisfies $\sum_{y\in{\mathcal{X}}}a_{y}=1$ and $a_{y}=0$ for all $y\notin S(x)$, we see that
@@
$\Phi_{x,z}(\alpha)=\underset{a\in\mathbb{R}^{d}}{\mathrm{arg\,max}}\,\Bigl{\{}\sum_{y\in S(x)}\!a_{y}\,(-\mathbb{E}[\varphi(y,Z^{\prime}(x,z,\alpha))])-f(a)\Bigr{\}}.$
@@
$|\Phi_{x,z}(\alpha)-\Phi_{x,z}(\tilde{\alpha})|_{1}\leq\frac{h}{\gamma}\sum_{y\in S(x)}\mathbb{E}|\varphi(y,Z^{\prime}(x,z,\alpha))-\varphi(y,Z^{\prime}(x,z,\tilde{\alpha}))|.$
@@
$\sum_{y\in S(x)}\mathbb{E}|\varphi(y,Z^{\prime}(x,z,\alpha))-\varphi(y,Z^{\prime}(x,z,\tilde{\alpha}))|$ $\leq\sum_{y\in S(x)}\mathrm{Lip}_{1}(\varphi(y,\cdot))\,\mathbb{E}|Z^{\prime}(x,z,\alpha)-Z^{\prime}(x,z,\tilde{\alpha})|_{1}$
@@
$\leq L_{\varphi}\,\max_{(x,z)\in\mathscr{X}_{N}}|\alpha(x,z)-\tilde{\alpha}(x,z)|_{1}$
$=L_{\varphi}\,\max_{(x,z)\in\mathscr{X}_{N}}\sum_{y\in S(x)}|\alpha_{y}(x,z)-\tilde{\alpha}_{y}(x,z)|,$
@@
$|\Phi(\alpha)-\Phi(\tilde{\alpha})|_{\mathcal{S}}\leq\frac{h}{\gamma}\max_{(x,z)\in\mathscr{X}_{N}}\sum_{y\in S(x)}\mathbb{E}|\varphi(y,Z^{\prime}(x,z,\alpha))-\varphi(y,Z^{\prime}(x,z,\tilde{\alpha}))|\leq\frac{L_{\varphi}h}{\gamma}|\alpha-\tilde{\alpha}|_{\mathcal{S}}.$
Hence, $\Phi(\cdot)$ is a contraction if $h<\gamma/L_{\varphi}$, in which case Banach’s fixed-point theorem guarantees the uniqueness of a fixed point $\hat{\alpha}\in{\mathcal{S}}$.
We proceed to establish an upper bound of the Lipschitz constant of $\hat{\alpha}(x,\cdot)$. Fix $z,\tilde{z}\in\Sigma^{d-1}_{N}$. By Assumption 2.7, for any $a,\tilde{a}\in A(h,x)$ and $\tilde{w}\in\partial_{a}\ell(x,\tilde{z},a/h)\subset\mathbb{R}^{d}$,
@@
$H(x,\tilde{z},$ $\,E(x,\tilde{z},\varphi,\hat{\alpha}),\tilde{a})=\ell(x,\tilde{z},\tilde{a}/h)h+\tilde{a}\cdot E(x,\tilde{z},\varphi,\hat{\alpha})$
@@
Now we apply this to $a=\hat{\alpha}(x,z)$ and $\tilde{a}=\hat{\alpha}(x,\tilde{z})$. By the definition of $\hat{\alpha}(x,\tilde{z})$ as a minimizer of $H(x,\tilde{z},E(x,\tilde{z},\varphi,\hat{\alpha}),\cdot\,;\hat{\alpha})$,
$H(x,\tilde{z},E(x,\tilde{z},\varphi,\hat{\alpha}),\hat{\alpha}(x,z))$ $\geq H(x,\tilde{z},E(x,\tilde{z},\varphi,\hat{\alpha}),\hat{\alpha}(x,\tilde{z}))$
@@
As $\hat{\alpha}(x,z)$ minimizes $H(x,z,E(x,z,\varphi,\hat{\alpha}),\,\cdot\,)$, by [7, Corollary 3.68] there exists a $w\in\partial_{a}\ell(x,z,\hat{\alpha}(x,z)/h)$ such that
@@
00 0 $\geq[\tilde{w}+E(x,\tilde{z},\varphi,\hat{\alpha})]\cdot(\hat{\alpha}(x,\tilde{z})-\hat{\alpha}(x,z))+\frac{\gamma}{h}|\hat{\alpha}(x,\tilde{z})-\hat{\alpha}(x,z)|_{1}\,|\hat{\alpha}(x,\tilde{z})-\hat{\alpha}(x,z)|_{\infty}$
@@
$\sum_{y\in S(x)}\mathbb{E}|\varphi(y,Z^{\prime}(x,\tilde{z},\hat{\alpha}))-\varphi(y,Z^{\prime}(x,z,\hat{\alpha}))|$ $\leq L_{\varphi}\ \mathbb{E}|Z^{\prime}(x,\tilde{z},\hat{\alpha})-Z^{\prime}(x,z,\hat{\alpha})|_{1}$
$\leq L_{\varphi}\ (1+\max_{x\in{\mathcal{X}}}\mathrm{Lip}_{1}(\hat{\alpha}(x,\cdot)))\,|z-\tilde{z}|_{1}.$
@@
$\frac{\gamma}{h}|\hat{\alpha}(x,\tilde{z})-\hat{\alpha}(x,z)|_{1}$ $\leq\sum_{y\in S(x)}\bigl{(}|\tilde{w}(y)-w(y)|+\mathbb{E}|\varphi(y,Z(x,\tilde{z},\hat{\alpha}))-\varphi(y,Z(x,z,\hat{\alpha}))|\bigr{)}$
$\leq\bigl{(}L_{\partial\ell}+L_{\varphi}(1+\max_{x\in{\mathcal{X}}}\,\mathrm{Lip}_{1}(\hat{\alpha}(x,\cdot)))\bigr{)}\,|z-\tilde{z}|_{1}.$
@@
Using Assumption 2.5 (iii), we know that there exists a $b_{*}<\infty$ such that the value function (2.5) of the tagged player is bounded by $b_{*}(1+T)$, uniformly in the step size $h\in(0,\bar{h})$ for any $\bar{h}<\infty$. Now note that any function $\varphi:\Sigma^{d-1}_{N}\to\mathbb{R}$ trivially satisfies
@@
Let Assumptions 2.5 and 2.7 hold. Let $h>0$ be a time step and $T\in h\mathbb{Z}_{+}$ an arbitrary time horizon. If
$h<h_{*}(N,T):=\frac{\gamma}{2m(d)Nb_{*}(1+T)},$
@@
[…]
@@
For a (sufficiently small) thermal noise level $\sigma^{2}\geq 0$ set
${\mathcal{X}}=\{0,1\},\quad A(h,x)=\{(a_{0},a_{1})\in\Sigma^{1}:a_{0},a_{1}\geq\sigma^{2}h\},\quad\ell(x,a)=\frac{1}{2h}a_{x+1}^{2},\quad g(x^{i},x^{-i})=\chi_{x^{i}\neq x^{-i}},$
for $(x^{0},x^{1},a)\in{\mathcal{X}}^{2}\times\Sigma^{1}$. Addition is understood modulo $2$, and we observe that the cost function $g$ encourages players to be in the same state at terminal time while $\ell$ penalizes changing one’s current state. In this two-state game, it is equivalent to identify controls $\alpha:{\mathcal{X}}\to\Sigma^{1}$ with the probability of changing, i.e., $\alpha(x)=\alpha_{x+1}(x)$. Then, admissible controls are given by vectors $(\alpha(0),\alpha(1))\in\tilde{A}(h)^{2}$ where $\tilde{A}(h):=[\sigma^{2}h,1-\sigma^{2}h]$. Given a control $\beta=(\beta(0),\beta(1))\in\tilde{A}(h)^{2}$ of the other player, the tagged player solves
$\inf_{\alpha:\,{\mathcal{X}}\to\tilde{A}(h)}\alpha(x)^{2}/(2h)+\mathbb{E}^{\alpha\otimes\beta}[g(X^{0}_{h},X^{1}_{h})\,|\,X_{0}=x],$
for every $x\in{\mathcal{X}}^{2}$. Here, $\mathbb{P}^{\alpha}[X^{0}_{h}=x^{0}+1\,|\,X^{0}_{0}=x^{0}]=\alpha(x^{0})$, $x^{0}\in{\mathcal{X}}$. A direct computation shows that the NLL equation characterizing symmetric MPE reduces to the following system of equations:
@@
and $\alpha(1,1)=\alpha(0,0)$. One easily observes the necessity for small $h$ to ensure the unique solvability of this set of equations. Indeed, if $\sigma^{2}=0$, then the first equation possesses multiple solutions for $h\geq 1$, two of them being $\alpha(0,0)=0,1$. If $\sigma^{2}>0$, another direct computation shows the existence of three solutions if $h\in(h_{*},h^{*})$ where $h_{*}:=(1+\sigma^{2}-\sqrt{1-(6-\sigma^{2})\sigma^{2}})/(4\sigma^{2})$ and $h^{*}:=(1+\sigma^{2}+\sqrt{1-(6-\sigma^{2})\sigma^{2}})/(4\sigma^{2})$. This assumes that $\sigma^{2}\leq 3-2\sqrt{2}$, so that the square root is well-defined.
@@ 4.1 The mean-field game system
We continue to establish the existence result in Proposition 2.18 (i).
@@
Set ${\mathcal{S}}:=(\Sigma^{d-1})^{{\mathcal{T}}(t,T)}$ and define a map $\Phi:{\mathcal{S}}\to{\mathcal{S}}$ via the following procedure. For $\boldsymbol{\mu}\in{\mathcal{S}}$, let $(v,\alpha)$ solve the dynamic programming equation
$\begin{cases}v(s,x)=\underset{a\in A(h,x)}{\inf}\ H(x,\mu_{s},v(s+h),a),\quad t\leq s\leq T,\\ \alpha(s,x)=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\ H(x,\mu_{s},v(s+h),a),\quad t\leq s<T,\\ v(T,x)=g(x,\mu_{T}).\end{cases}$
@@
With this construction, it is clear that solutions to the MFG system correspond to fixed points of the map $\Phi(\cdot)$, and hence by Remark 2.15, correspond to mean-field game Nash equilibria. But Berge’s maximum theorem implies that $\Phi(\cdot)$ is a continuous map. The claim follows by Brouwer’s fixed-point theorem. ∎

### v2

@@ 3.1 Existence
$\begin{cases}v^{\beta}(t,x,z)=\underset{a\in A(h,x)}{\inf}\,H(x,z,E(x,z,v^{\beta}(t+h),\beta(t)),a),\\ \alpha(t,x,z)=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\ H(x,z,E(x,z,v^{\beta}(t+h),\beta(t)),a),\\ v^{\beta}(T,x,z)=g(x,z),\end{cases}$
by backward recursion. Then set $\Phi(\beta):=\alpha$. Using Berge’s maximum theorem, it is easy to see that $\Phi$ is a continuous mapping. Since $\mathscr{A}_{N}$ is a convex compact set, Brouwer’s fixed-point theorem guarantees the existence of a fixed point. By Remark 2.7, any fixed point corresponds to a symmetric Markov perfect equilibrium. ∎
@@ 3.2 Uniqueness
$\eta_{j}\overset{d}{=}\sum_{i=1}^{n}\chi_{[0,p_{j}]}(\xi^{i})\in\mathbb{Z}_{+},$
where $\{\xi^{i}\}$ are independent $\mathrm{Uniform}(0,1)$ random variables. Now fix $j\in{\mathcal{X}}$ and $(x,z)\in\mathscr{X}_{N}$, let $\xi^{i}_{y}$, $1\leq i\leq N,\,y\in{\mathcal{X}}$, be a collection of independent $\mathrm{Uniform}(0,1)$ random variables, and write
@@
$\overset{d}{=}\frac{1}{N}\sum_{y\in{\mathcal{X}}}\sum_{i=1}^{Nz_{y}}(\chi_{[0,\alpha_{j}(y,z+e^{N}_{xy})]}-\chi_{[0,\tilde{\alpha}_{j}(y,z+e^{N}_{xy})]})(\xi^{i}_{y}).$
Taking expectations,
@@
$\leq\,(1+\max_{y\in{\mathcal{X}}}\mathrm{Lip}_{1}(\alpha(y,\cdot)))\,|z-\tilde{z}|_{1}.$
Taking the maximum over $(x,z)$ again finishes the proof. ∎
We note the following lemma, whose proof is given in Appendix A.2.
@@
Let $f:\mathbb{R}^{d}\to(-\infty,\infty]$ be a proper, closed and convex function. Let $C\subset\mathrm{dom}(f)$ be a non-empty, closed and convex set. Fix $p,q\in[1,\infty]$ with $1/p+1/q=1$ and assume that there exists a $\gamma>0$ such that
$f(x^{\prime})\geq f(x)+\phi\cdot(x^{\prime}-x)+\frac{\gamma}{2}|x^{\prime}-x|_{p}|x^{\prime}-x|_{q}$ (3.1)
for any $x,x^{\prime}\in C$ and $\phi\in\partial f(x)$. Set $\psi(y)=\mathrm{arg\,max}\{y\cdot x-f(x):x\in C\}$, $y\in\mathbb{R}^{d}$. Then, $\psi$ is single-valued and is Lipschitz continuous with constant $1/\gamma$, with respect to both $|\cdot|_{p}$ and $|\cdot|_{q}$.
We are now ready to prove existence and uniqueness for one-step games. We recall the constant $\gamma$ from Assumption 2.8, $S(x)$ and $m$ from Definition 2.2, and the constant $L_{\partial\ell}$ from Assumption 2.20.
@@
Let Assumptions 2.6 and 2.8 hold. Then, for any $\varphi:\mathscr{X}_{N}\to\mathbb{R}$, there exists a unique function $\hat{\alpha}:\mathscr{X}_{N}\to\Sigma^{d-1}$ that satisfies
@@
For $\alpha,\tilde{\alpha}\in{\mathcal{S}}$ set $d_{\mathcal{S}}(\alpha,\tilde{\alpha}):=\max_{(x,z)\in\mathscr{X}_{N}}|\alpha(x,z)-\tilde{\alpha}(x,z)|_{1}$,
@@
Now let $(x,z)\in\mathscr{X}_{N}$ be given. Define a function $f:\mathbb{R}^{d}\to[0,\infty]$ by $f(a):=\ell(x,z,a/h)h$ for $a\in[0,\infty)^{d}$ and $\infty$ otherwise. Set $C:=A(h,x)$. By Assumptions 2.6 and 2.8, $f$ is a proper, closed and convex function. Using Assumption 2.8 and the fact that any element $a\in C$ satisfies $\sum_{y\in{\mathcal{X}}}a_{y}=1$ and $a_{y}=0$ for all $y\notin S(x)$, we see that
@@
$\Phi_{x,z}(\alpha)=\underset{a\in C}{\mathrm{arg\,max}}\,\Bigl\{\sum_{y\in S(x)}\!a_{y}\,(-\mathbb{E}[\varphi(y,Z^{\prime}(x,z,\alpha))])-f(a)\Bigr\}.$
@@
[…]
@@
[…]
@@
$\leq L_{\varphi}\,\max_{(x,z)\in\mathscr{X}_{N}}|\alpha(x,z)-\tilde{\alpha}(x,z)|_{1},$
@@
[…]
Hence, $\Phi$ is a contraction since $h<\gamma/L_{\varphi}$, in which case Banach’s fixed-point theorem guarantees the uniqueness of a fixed point $\hat{\alpha}\in{\mathcal{S}}$.
We proceed to establish an upper bound of the Lipschitz constant for $\hat{\alpha}(x,\cdot)$. Fix $z,\tilde{z}\in\Sigma^{d-1}_{N}$. By Assumption 2.8, for any $a,\tilde{a}\in A(h,x)$ and $\tilde{w}\in\partial_{a}\ell(x,\tilde{z},a/h)\subset\mathbb{R}^{d}$,
@@
[…]
@@
[…]
$H(x,\tilde{z},E(x,\tilde{z},\varphi,\hat{\alpha}),\hat{\alpha}(x,z))$ $)\displaystyle\geq H(x,\tilde{z},E(x,\tilde{z},\varphi,\hat{\alpha}),\hat{\alpha}(x,\tilde{z}$))
@@
As $\hat{\alpha}(x,z)$ minimizes $)H(x,z,E(x,z,\varphi,\hat{\alpha}$$\cdot\,),$ by [7, Corollary 3.68] there exists a $w\in\partial_{a}\ell(x,z,\hat{\alpha}(x,z)/h)$ such that
@@
$0$ $\geq[\tilde{w}+E(x,\tilde{z},\varphi,\hat{\alpha})]\cdot(\hat{\alpha}(x,\tilde{z})-\hat{\alpha}(x,z))+\frac{\gamma}{h}|\hat{\alpha}(x,\tilde{z})-\hat{\alpha}(x,z)|_{1}\,|\hat{\alpha}(x,\tilde{z})-\hat{\alpha}(x,z)|_{\infty}$
@@
[…]
$\leq L_{\varphi}\ (1+\max_{x\in{\mathcal{X}}}\mathrm{Lip}_{1}(\hat{\alpha}(x,\cdot)))\,|\tilde{z}-z|_{1}.$
@@
[…]
$\leq\bigl(L_{\partial\ell}+L_{\varphi}(1+\max_{x\in{\mathcal{X}}}\,\mathrm{Lip}_{1}(\hat{\alpha}(x,\cdot)))\bigr)\,|\tilde{z}-z|_{1}.$
@@
Using Assumption 2.6 (iv), we know that there exists a $b_{*}<\infty$ such that the value function (2.5) of the tagged player is bounded by $b_{*}(1+T)$, uniformly in the step size $h\in(0,\bar{h})$ for any $\bar{h}<\infty$. Now note that any function $\varphi:\Sigma^{d-1}_{N}\to\mathbb{R}$ trivially satisfies
@@
Let Assumptions 2.6 and 2.8 hold. Let $h>0$ be a time step and $T\in h\mathbb{Z}_{+}$ an arbitrary time horizon. If
$h<h_{*}(N,T):=\frac{\gamma}{2mNb_{*}(1+T)},$
@@
The following example of a two-player game illustrates the necessity of the small-$h$ condition to ensure the existence and uniqueness of a MPE. We will analyze the mean-field version of this example in Subsection 4.3.
@@
For $N=1$ and a (sufficiently small) thermal noise level $\sigma^{2}\geq 0$, set
${\mathcal{X}}=\{0,1\},\quad A(h,x)=\{(a_{0},a_{1})\in\Sigma^{1}:a_{0},a_{1}\geq\sigma^{2}h\},\quad\ell(x,a)=\frac{1}{2}a_{x+1}^{2},\quad g(x,z)=\chi_{x\neq z},$
for $(x,z,a)\in{\mathcal{X}}^{2}\times\Sigma^{1}$. Addition is understood modulo $2$, and we observe that the cost function $g$ encourages players to be in the same state at terminal time while $\ell$ penalizes changing one’s current state. In this two-state game, we can identify controls $\alpha:{\mathcal{X}}^{2}\to\Sigma^{1}$ with the probability of changing, i.e., write $\alpha(x,z)=\alpha_{x+1}(x,z)$. Then, admissible controls are vectors $\alpha(x,z)\in\tilde{A}(h)=:[\sigma^{2}h,1-\sigma^{2}h]$. Given a control $\beta(x,z)$ of the other player, the tagged player solves
$\inf_{\alpha:\,{\mathcal{X}}^{2}\to\tilde{A}(h)}\alpha(x,z)^{2}/(2h)+\mathbb{E}^{\alpha\otimes\beta}[g(X_{h},Z_{h})\,|\,(X_{0},Z_{0})=(x,z)],$
for every $(x,z)\in{\mathcal{X}}^{2}$. A direct computation shows that the NLL equation characterizing symmetric MPE reduces to the following system of equations:
@@
and $\alpha(1,1)=\alpha(0,0)$. One easily observes the necessity for small $h$ to ensure the unique solvability of this set of equations. Indeed, if $\sigma^{2}=0$, then the first equation possesses multiple solutions for $h\geq 1$, two of them being $\alpha(0,0)\in\{0,1\}$. If $\sigma^{2}>0$, another direct computation shows the existence of three solutions if $h\in(h_{*},h^{*})$ where $h_{*}:=(1+\sigma^{2}-\sqrt{1-(6-\sigma^{2})\sigma^{2}})/(4\sigma^{2})$ and $h^{*}:=(1+\sigma^{2}+\sqrt{1-(6-\sigma^{2})\sigma^{2}})/(4\sigma^{2})$. This assumes that $\sigma^{2}\leq 3-2\sqrt{2}$, so that the square root is well-defined.
@@ 4.1 The mean-field game system
We continue to establish the existence result in Proposition 2.19 (i). The proof of 2.19 (ii) follows by inspection and hence we omit it.
@@
Set ${\mathcal{S}}:=(\Sigma^{d-1})^{{\mathcal{T}}(t,T)}$ and define a map $\Phi:{\mathcal{S}}\to{\mathcal{S}}$ via the following procedure. For $\boldsymbol{\mu}\in{\mathcal{S}}$, let $(v,\alpha)$ solve the dynamic programming equation for $t\leq s<T$:
$\begin{cases}v(s,x)=\underset{a\in A(h,x)}{\inf}\ H(x,\mu_{s},v(s+h),a),\\ \alpha(s,x)=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\ H(x,\mu_{s},v(s+h),a),\\ v(T,x)=g(x,\mu_{T}).\end{cases}$
@@
With this construction, it is clear that solutions to the MFG system correspond to fixed points of the map $\Phi$, and hence by Remark 2.16, correspond to mean-field game Nash equilibria. Berge’s maximum theorem implies that $\Phi$ is a continuous map. The claim follows by Brouwer’s fixed-point theorem. ∎